OBSERVATIONS OF INTERSTELLAR C_{3} IN TRANSLUCENT SIGHTLINES

\underline{T. OKA}, J. A. THORBURN, L. M. HOBBS, D. E. WELTY, D. G. YORK, Department of Astronomy & Astrophysics, The University of Chicago, Chicago, IL 60637; B. J. MCCALL, Department of Astronomy, University of California, Berkeley, CA 94720; S. D. FRIEDMAN and P. SON-NENTRUCKER, Department of Physics and Astronomy, Johns Hopkins University, Baltimore, MD 21218.

Following the detection of the 000-000 band of the A^{1}Π_{u} ← X^{1}Σ_{g}^{+} transition of C_{3} at 4051.6 Å by Maier and others^{a} toward three diffuse sightlines, we have searched for the spectrum toward more reddened stars. Although the lower resolution (R = 37,500) of our spectrometer (ARCES) did not resolve the rotational structure of the spectrum, it allowed us to observe denser clouds toward fainter stars which are expected to contain more carbon molecules. The C_{3} spectrum has been observed toward 15 stars with color excesses ranging from E(B-V) = 0.33 to 1.12 and the C_{3} column densities were determined by comparing the observed spectra with simulated spectra. The observed C_{3} column densities from 10^{12} to 10^{13} cm^{−2} are well correlated with the C_{2} column densities indicating their close chemical relation. The formation of C_{3} from C_{2} will be discussed. The studies of C_{3} and C_{2} have led us to the revelation of a group of diffuse interstellar bands whose intensities correlate well with the C_{3}/C_{2} column densities^{b}. It was also noted that abundances of C_{3}/C_{2} and H_{3}^{+} are completely uncorrelated.

Our non-detection of the 020-000 vibronic band at 3991.6 Å set an upper limit to its Franck-Condon factor and that of C_{4} and C_{5} set upper limits of their column densities.

aJ. P. Maier, N. M. Lakin, G. M. Walker and D. A. Bohlender Astrophys. J. \underline{553}, 267, 2001

bB. J. McCall et al., this meeting